OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10 -2y + 10 2y + 10 2y - 10 -2y - 10

OpenStudy (nurali):

3x + y = 8----(1) -x - 2y = -10---(2) in equation (2) -x=-10+2y x=-(-10+2y) x=10-2y

OpenStudy (anonymous):

alright thanks good night

OpenStudy (nurali):

good night.